Help The following frequency distribution shows the ages of Indias 62 richest individuals. One of these individuals is selected at random. es Frequency 30up to 40 40 up to 50 50 up to 60 60, up to 70 70 up to 80 13 14 14 10 a. What is the probability that the individual is between 50 and 60 years of age? (Round your answer to 3 decimal places.) Probability years of age? (Round your answer to 3 decimal places) b. What is the probability that the individual is younger than 5 c. What is the probability that the individual is at least 60 years of age? (Round your answer to 3 decimal places.) Probability
Answer #1

Total frequency = 62

a. Probability = 14/62 = 0.226

b. Probability = (11+13)/62 = 0.387

c. Probability = (14+10)/62 = 0.387